OpenStudy (anonymous):

Solve for x using logs. 8 * 5^x = 7 * 3^x x=________. How do I solve this using logs? :/

OpenStudy (agent0smith):

Start by taking logs of both sides.

OpenStudy (anonymous):

so log(8 * 5^x)= log(7*3^x) ? or is it just log(8) * 5^x = log(7) * 3^x ?

OpenStudy (dumbcow):

the first option then split it up like the previous problem remember \[\ln a^x = x \ln a\]

OpenStudy (anonymous):

oh okay :) sorry what does that last part say? latex still isn't showing up :(

OpenStudy (dumbcow):

oh log(a^x) = xlog(a)

OpenStudy (anonymous):

so log(8) x log(5^x)=log(7) x log (3^x) is that how you split it up?

OpenStudy (agent0smith):

log(ab) = loga + logb so log8 + log 5^x = log7 + log3^x

OpenStudy (anonymous):

ohh okay so would it become this? 5^x log (8) = 3^x log (7) ?

OpenStudy (anonymous):

oh so i don't need the parentheses?

OpenStudy (agent0smith):

No, log(8 * 5^x)= log(7*3^x) becomes log8 + log 5^x = log7 + log3^x

OpenStudy (anonymous):

okay so it simplifies to this? .903 + log 5^x = .845 + log 3^x ?? not really sure how to simplify the logs with the x as the exponent though :/

OpenStudy (dumbcow):

use the property i posted ^^^

OpenStudy (anonymous):

okay so log(a^x) = xlog(a) that makes it .903 + xlog(5) = .845 + xlog(3) ?

OpenStudy (dumbcow):

:)

OpenStudy (anonymous):

.903 + x(.699) = .845 + x(.477) is that right? :/

OpenStudy (dumbcow):

yeah, by changing the logs to rounded decimals it may effect the precision of your final answer just a warning

OpenStudy (anonymous):

ohh okay here are the exact numbers .903089987 + x (.6989700043) = .84509804 + x (.4771212547) 0.057991947 + x(.6989700043)=x (.4771212547) is that right so far?

OpenStudy (dumbcow):

oh lol you were right before, this is no different fyi the exact numbers go on forever (they are irrational) anyway go ahead and solve for x by isolating it on one side

OpenStudy (anonymous):

haha okay :P to isolate, what do i do? :/ am I dividing one of them?

OpenStudy (dumbcow):

oh sorry i missed you already started by combining the constants (correct) move the .69x to other side by subtracting

OpenStudy (anonymous):

0.057991947 + x=x (-0.2218487496) ??

OpenStudy (dumbcow):

.47x - .69x = -.22x get rid of the extra "+x" on left side

OpenStudy (anonymous):

ohh okay.. what happens next?

OpenStudy (dumbcow):

how do you solve this 7x = 21 ??

OpenStudy (anonymous):

divide by 7 x=3 ?

OpenStudy (dumbcow):

right :) now apply that to your problem -.22x = .05

OpenStudy (anonymous):

wait so does x=-4.104954486 ? not sure if that's right? :/

OpenStudy (anonymous):

oh wait... is it -.227 ?? so about -0.23 ?

OpenStudy (anonymous):

is that right? x=-0.23 ?

OpenStudy (anonymous):

oh so the solution is actually x=-1.567 ?

OpenStudy (dumbcow):

x = -.2614

OpenStudy (anonymous):

ohh okay

OpenStudy (anonymous):

how did you get that? :/ i think i did the math wrong haha

OpenStudy (dumbcow):

you did it right but you had a rounding error use the exact numbers right from calculator to avoid rounding error

OpenStudy (anonymous):

ohh okay cool! will keep that in mind :) thanks so much!!! :D